OpenStudy (anonymous):

What is the inverse of log (base 3) 2x?

OpenStudy (watchmath):

Let \(y=\log_3(2x)\) Then \(3^y=2x\) \(x=\frac{1}{2}\cdot 3^y\) Switch between \(x\) and \(y\) we have \(y=\frac{1}{2}\cdot 3^x\) as the inverse function.